# Full spectrum of Love numbers of Reissner-Nordström black hole in $D$-dimensions

### v2

@@ ABSTRACT
xiaminghao@tju.edu.cn
maliang0@tju.edu.cn
@@ 1 Introduction
Tidal Love numbers describe the deformability of a black hole under the influence of external force. It has been demonstrated that conservative linear tidal Love numbers vanish for several families of asymptotically flat black holes in four dimensions, including Schwarzschild [1], Kerr [2, 3, 4] and Reissner-Nordström (RN) [5, 6, 7] black holes. It was also found that for Kerr black holes, the conservative tidal Love numbers remain vanishing at nonlinear level [8, 9, 10, 11]. These results provide strong support to the no-hair theorem, showing that black holes in four-dimensional general relativity are remarkably rigid objects. On the other hand, Love numbers are generically nonvanishing for charged black holes coupled to charged scalar or fermion [12, 13, 14, 15] and black holes in higher dimensions [16, 17, 18, 19], in modified gravitational theories [20, 21, 22, 23, 24, 25, 26], in the presence of a cosmological constant [27, 28], or in astrophysical environments [29, 30, 31, 32, 33]. Its potential observability by future gravitational wave experiments makes Love numbers useful theoretical diagnostics.
@@
We carry out the analysis by expanding the $D$-dimensional fields in terms of harmonic functions on $S^{D-2}$, plugging them into the Einstein-Maxwell action, linearized about the background RN black hole to yield an effective 2-dimensional quadratic action for perturbative fields in $1+1$ dimensions. Upon diagonalizing the perturbations, one obtains the complete set of master equations generalizing previous results for Schwarzschild black hole [38, 39, 40, 41, 42]. Although the tensor and vector modes were previously analyzed using the Ishibashi-Kodama formalism [6], we rederive them here within the effective-action framework to ensure a unified and self-consistent treatment of all perturbations channels, which is also necessary for cross-checking known results.
@@ 2 Perturbations about the RN black hole
$\bar{R}_{\mu\nu}-\frac{1}{2}\bar{g}_{\mu\nu}\bar{R}-\frac{1}{2}\bar{F}_{\mu\nu}^{2}+\frac{1}{8}\bar{g}_{\mu\nu}\bar{F}^{2}=0,\quad\nabla_{\mu}\bar{F}^{\mu\nu}=0\ .$ (3)
@@
$d\bar{s}_{D}^{2}$ $=$ $-fdt^{2}+\frac{dr^{2}}{f}+r^{2}d\Omega^{2}_{D-2}\,,\qquad f=1-\frac{2\mu}{r^{D-3}}+\frac{q^{2}}{r^{2(D-3)}}\ ,$ (9)
$\bar{A}_{(1)}$ $=$ $\psi dt\,,\qquad\psi=\sqrt{\frac{2(D-2)}{D-3}}\frac{q}{r^{D-3}}\ .$ (10)
@@
$\qquad\qquad\qquad+\frac{(\ell+1)(\ell+D-4)-2(D-3)f-2rf^{\prime}}{r^{2}f}h_{0}^{2}\Bigg]\ ,$
@@
$\qquad\quad+H_{0}\Bigg(\big(f^{\prime}+\frac{2(D-3)f}{r}\big)\mathcal{H}_{1}+2f\mathcal{H}_{1}^{\prime}-\Big(1+\frac{(D-2)\big((D-3)f+rf^{\prime}\big)}{\ell(\ell+D-3)}\Big)H_{2}$
$\qquad\quad-\frac{(D-2)rfH_{2}^{\prime}}{\ell(\ell+D-3)}\Bigg)-\frac{2(D-3)f+rf^{\prime}}{r}H_{2}\mathcal{H}_{1}+H_{1}\Big(H_{1}-2\dot{\mathcal{H}}_{1}+\frac{2(D-2)r}{\ell(\ell+D-3)}\dot{H}_{2}\Big)\Bigg]$
$\qquad\quad+\frac{r^{D-3}}{8}\Big(rf^{\prime\prime}+(D-2)f^{\prime}\Big)\Big((H_{0}+H_{2})^{2}-4H_{1}^{2}\Big)\ .$ (27)
@@ 3.1 Tensor Love numbers in diverse dimensions
where the last term encodes the contribution from the background electric field. To rewrite the action in standard form, we redefine the field $\Psi_{T}$
@@
    $R=\frac{(-1)^{2\tilde{\ell}}\Gamma(\tilde{\ell}+1)^{2}}{(2\tilde{\ell})!(2\tilde{\ell}+1)!\Gamma(-\tilde{\ell})^{2}}(1-\sigma)^{2\tilde{\ell}+1},\quad\sigma:=\tilde{r}_{-}/\tilde{r}_{+}\$. (41)
    When $\tilde{\ell}$ is an integer, the coefficient $R$ vanish. Meanwhile, the solution regular at the horizon consists of only $\Phi_{\rm tidal}$ which is a polynomial of degree $2\tilde{\ell}+1$. Therefore, when $\tilde{\ell}$ is an integer, tensor Love number always vanishes.
@@
    $k_{T}=\frac{\Gamma(-2\tilde{\ell}-1)\Gamma(\tilde{\ell}+1)^{2}}{\Gamma(2\tilde{\ell}+1)\Gamma(-\tilde{\ell})^{2}}(1-\sigma)^{2\tilde{\ell}+1}\quad\text{(in terms of $z$)}\ ,$ (42)
@@ 3.2 Effective action of vector type perturbations
$f^{\prime\prime}+\frac{D-2}{r}f^{\prime}-\frac{D-3}{D-2}\psi^{\prime 2}=0,\quad\psi^{\prime\prime}+\frac{D-2}{r}\psi^{\prime}=0\ .$ (48)
@@
$\alpha_{1}=1,\quad\alpha_{2}=\alpha_{3}=0,\quad\alpha_{4}=-\sqrt{(\ell-1)(\ell+D-2)}.$ (63)
@@ 3.3 Vector-type Love numbers in diverse dimensions
$\alpha$ $=$ $-\frac{D-2}{D-3}-\tilde{\ell},\quad\beta=\tilde{\ell}-\frac{1}{D-3}\ .$ (67)
@@ 3.4 Effective action of scalar type perturbations
$a_{2}$ $=$ $\frac{c_{2}r^{D-3}+(2-D)q}{r^{D-3}\sqrt{c_{2}^{2}+(\ell-1)(\ell+D-2)}},\quad a_{4}=\frac{\sqrt{\mathcal{F}}}{\sqrt{2(D-3)}\sqrt{c_{2}^{2}+(\ell-1)(\ell+D-2)}}\ ,$ (90)
@@
where the potentials in equations above are postponed to the appendix (A) and (120) . For $D=4$, we find that the equations of motion for $\mathcal{U}_{\pm}$ agree with those in [44] where the corresponding functions are denoted by $R_{\pm}$. For Schwarzschild black hole ($q=0$) the coefficients in (90) reduce to
@@ 4 Conclusions
In this work, we focused on the purely electric RN black hole. It is conceivable that via electromagnetic duality, our results can be generalized to dyonically charged black hole. We would like to consider spherically symmetric black hole in Einstein-Maxwell-dilaton system, some of which do admit analytically solutions. Finally, it should be interesting to consider nonlinear tidal Love numbers for charged black holes as was done for the neutral ones [8].
@@ Acknowledgement
Appendix A Explicit forms of the potentials in the scalar sector

### v3

@@ ABSTRACT
Keywords— Higher dimensional black holes, Gravity in more than four dimensions, Physics of black holes, Classical black holes
@@ 1 Introduction
Tidal Love numbers describe the deformability of a black hole under the influence of external force. It has been demonstrated that conservative linear tidal Love numbers vanish for several families of asymptotically flat black holes in four dimensions, including Schwarzschild [1], Kerr [2, 3, 4] and Reissner-Nordström (RN) [5, 6, 7] black holes. It was also found that for Kerr black holes, the conservative tidal Love numbers remain vanishing at nonlinear level [8, 9, 10, 11]. These results provide strong support to the no-hair theorem, showing that black holes in four-dimensional general relativity are remarkably rigid objects. On the other hand, Love numbers are generically nonvanishing for dynamical black holes [12, 13, 14], charged black holes coupled to charged scalar or fermion [15, 16, 17, 18] and black holes in higher dimensions [19, 20, 21, 22], in modified gravitational theories [23, 24, 25, 26, 27, 28, 29], in the presence of a cosmological constant [31, 32], or in astrophysical environments [33, 34, 35, 36, 37]. Therefore, Love numbers are useful theoretical diagnostics. In fundamental theories, such as string theory or supergravit theory, Love numbers might be used to distinguish black holes from other compact objects, such as the black rings [38] or fuzzball solutions [39]. Potential detectability by future gravitational wave experiments make Love numbers even more intriguing physical observables.
@@
We carry out the analysis by expanding the $D$-dimensional fields in terms of harmonic functions on $S^{D-2}$, plugging them into the Einstein-Maxwell action, linearized about the background RN black hole to yield an effective 2-dimensional quadratic action for perturbative fields in $1+1$ dimensions. Upon diagonalizing the perturbations, one obtains the complete set of master equations generalizing previous results for Schwarzschild black hole [44, 45, 46, 47, 48]. Although the tensor and vector modes were previously analyzed using the Ishibashi-Kodama formalism [6], we rederive them here within the effective-action framework to ensure a unified and self-consistent treatment of all perturbations channels, which is also necessary for cross-checking known results. Indeed, our results for Love numbers in tensor and vector sectors agree with previous findings in [6].
@@ 2 Perturbations about the RN black hole
$\bar{R}_{\mu\nu}-\frac{1}{2}\bar{g}_{\mu\nu}\bar{R}-\frac{1}{2}\bar{F}_{\mu\nu}^{2}+\frac{1}{8}\bar{g}_{\mu\nu}\bar{F}^{2}=0,\qquad\nabla_{\mu}\bar{F}^{\mu\nu}=0\ .$ (3)
@@
$d\bar{s}_{D}^{2}$ $=$ $-fdt^{2}+\frac{dr^{2}}{f}+r^{2}d\Omega^{2}_{D-2}\,,\hskip 18.49988ptf=1-\frac{2\mu}{r^{D-3}}+\frac{q^{2}}{r^{2(D-3)}}\ ,$ (9)
$\bar{A}_{(1)}$ $=$ $\psi dt\,,\hskip 18.49988pt\psi=\sqrt{\frac{2(D-2)}{D-3}}\frac{q}{r^{D-3}}\ .$ (10)
@@
$\hskip 18.49988pt\hskip 18.49988pt\hskip 18.49988pt+\frac{(\ell+1)(\ell+D-4)-2(D-3)f-2rf^{\prime}}{r^{2}f}h_{0}^{2}\Bigg]\ ,$
@@
$\hskip 18.49988pt\qquad+H_{0}\Bigg(\big(f^{\prime}+\frac{2(D-3)f}{r}\big)\mathcal{H}_{1}+2f\mathcal{H}_{1}^{\prime}-\Big(1+\frac{(D-2)\big((D-3)f+rf^{\prime}\big)}{\ell(\ell+D-3)}\Big)H_{2}$
$\hskip 18.49988pt\qquad-\frac{(D-2)rfH_{2}^{\prime}}{\ell(\ell+D-3)}\Bigg)-\frac{2(D-3)f+rf^{\prime}}{r}H_{2}\mathcal{H}_{1}+H_{1}\Big(H_{1}-2\dot{\mathcal{H}}_{1}+\frac{2(D-2)r}{\ell(\ell+D-3)}\dot{H}_{2}\Big)\Bigg]$
$\hskip 18.49988pt\qquad+\frac{r^{D-3}}{8}\Big(rf^{\prime\prime}+(D-2)f^{\prime}\Big)\Big((H_{0}+H_{2})^{2}-4H_{1}^{2}\Big)\ .$ (27)
@@ 3.1 Tensor Love numbers in diverse dimensions
where the last term comes from the energy momentum tensor of the background electric field, and thus vanishes in the case of Schwarzschild black hole. To rewrite the action in standard form, we redefine the field $\Psi_{T}$
@@
    $R=\frac{(-1)^{2\tilde{\ell}}\Gamma(\tilde{\ell}+1)^{2}}{(2\tilde{\ell})!(2\tilde{\ell}+1)!\Gamma(-\tilde{\ell})^{2}}(1-\sigma)^{2\tilde{\ell}+1},\qquad\sigma:=\tilde{r}_{-}/\tilde{r}_{+}\$. (41)
    Here and below,the effect of electric charge is encoded in the parameter $\sigma$. When $\tilde{\ell}$ is an integer, the coefficient $R$ vanish. Meanwhile, the solution regular at the horizon consists of only $\Phi_{\rm tidal}$ which is a polynomial of degree $2\tilde{\ell}+1$. Therefore, when $\tilde{\ell}$ is an integer, tensor Love number always vanishes.
@@
    $k_{T}=\frac{\Gamma(-2\tilde{\ell}-1)\Gamma(\tilde{\ell}+1)^{2}}{\Gamma(2\tilde{\ell}+1)\Gamma(-\tilde{\ell})^{2}}(1-\sigma)^{2\tilde{\ell}+1}\qquad\text{(in terms of $z$)}\ ,$ (42)
@@ 3.2 Effective action of vector type perturbations
$f^{\prime\prime}+\frac{D-2}{r}f^{\prime}-\frac{D-3}{D-2}\psi^{\prime 2}=0,\qquad\psi^{\prime\prime}+\frac{D-2}{r}\psi^{\prime}=0\ .$ (48)
@@
$\alpha_{1}=1,\qquad\alpha_{2}=\alpha_{3}=0,\qquad\alpha_{4}=-\sqrt{(\ell-1)(\ell+D-2)}.$ (63)
@@ 3.3 Vector-type Love numbers in diverse dimensions
$\alpha$ $=$ $-\frac{D-2}{D-3}-\tilde{\ell},\qquad\beta=\tilde{\ell}-\frac{1}{D-3}\ .$ (67)
@@ 3.4 Effective action of scalar type perturbations
$a_{2}$ $=$ $\frac{c_{2}r^{D-3}+(2-D)q}{r^{D-3}\sqrt{c_{2}^{2}+(\ell-1)(\ell+D-2)}},\qquad a_{4}=\frac{\sqrt{\mathcal{F}}}{\sqrt{2(D-3)}\sqrt{c_{2}^{2}+(\ell-1)(\ell+D-2)}}\ ,$ (90)
@@
where the potentials in equations above are postponed to the appendix (Appendix Explicit forms of the potentials in the scalar sector) and (120) . For $D=4$, we find that the equations of motion for $\mathcal{U}_{\pm}$ agree with those in [50] where the corresponding functions are denoted by $R_{\pm}$. For Schwarzschild black hole ($q=0$) the coefficients in (90) reduce to
@@ 4 Conclusions
In this work, we focused on the purely electric RN black hole. It is conceivable that via electromagnetic duality, our results can be generalized to dyonically charged black hole. We would like to consider spherically symmetric black hole in Einstein-Maxwell-dilaton system, some of which do admit analytically solutions. It should also be interesting to consider nonlinear tidal Love numbers for charged black holes as was done for the neutral ones [8]. In [58], the Love numbers of four-dimensional rotating STU black holes were analyzed. These results show that, for black holes in string theory and supergravity, the behavior of the Love numbers at the two-derivative level matches that of conventional black holes (e.g., Schwarzschild black holes). However, when higher-derivative terms are included, the strngy parameter $\alpha^{\prime}$ will also contributes to the Love numbers and potentially lift up the degenracy. Thus we would also like to investigate Love numbers of black holes in string theory with $\alpha^{\prime}$ corrections [59, 60, 62, 61]. Finally, our techinques may be generalized to study linear response of charged AdS black branes [64] or holes, which may lead to interesting applications in holography in diverse dimensions.
@@ Acknowledgement
Conflict of Interest   The authors declare that they have no conflict of interest.
Appendix Explicit forms of the potentials in the scalar sector